OpenStudy (anonymous):

If I flip four coins what is the probability that at least two of the coins are heads?

OpenStudy (anonymous):

Please help

OpenStudy (amistre64):

its the opposite of P(at most 1 is heads)

OpenStudy (amistre64):

1 - P(0) - P(1)

OpenStudy (anonymous):

do you have choices

OpenStudy (amistre64):

there are 16 outcomes tttt ttth ttht thtt httt ... 1 4 6 4 1 = 16 0 heads is 1 out of 16 1 head is 4 out of 16

OpenStudy (anonymous):

What about at least 2 heads

OpenStudy (amistre64):

at least 2 heads means that you have 2, 3, or 4 heads P(0) + P(1) + P(2) + P(3) + P(4) = 1 P(2) + P(3) + P(4) = 1 - P(0) - P(1)

OpenStudy (amistre64):

since P(0) = 1/16 and P(1) = 4/16 the solution is determined

OpenStudy (anonymous):

So what would the answer be written as?

OpenStudy (amistre64):

im not going to do the subtraction for you ....

OpenStudy (amistre64):

you can tell me what you get ... and i can verify if you did it correctly

OpenStudy (anonymous):

bro your even confusing me.... amistre64

OpenStudy (amistre64):

im not a mind reader, so youll have to be more detailed in what is confusing :)

OpenStudy (anonymous):

nevermind .. the probability would be 50% John

OpenStudy (amistre64):

the probability is not 50%; for starters, a probability is a number between 0 and 1 ... not a percentage.

OpenStudy (amistre64):

and for another, 1 - P(0) - P(1) does not equal .5000 or 1/2

OpenStudy (anonymous):

he asked for a probability. if you flip 4 coins and the odds of 2 of the comming out heads is 2/4 and tht equals 50% and wtf is p

OpenStudy (anonymous):

its not rocket science he just wants a simple answer.. idk whats wrong with tht

OpenStudy (amistre64):

\[\binom{4}{0}.5^0.5^4+\binom{4}{1}.5^1.5^3+\binom{4}{2}.5^2.5^2+\binom{4}{3}.5^3.5^1+\binom{4}{4}.5^n.4^0=1\] \[(1).5^0.5^4+(4).5^1.5^3+(6).5^2.5^2+(4).5^3.5^1+(1).5^4.5^0=1\]

OpenStudy (amistre64):

P(E) is notation for "probability of an event"

OpenStudy (amistre64):

the last 3 terms are 4 choose 2, 4 choose 3, and 4 choose 4 which is what we want if we are to determine "at least 2" out of 4

OpenStudy (amistre64):

subtracting the first 2 terms from each side gives us the solution

OpenStudy (anonymous):

your making this way more complex than it already isnt lol if you know the damn answer then tell him and quit arguing with a degenerate!

OpenStudy (amistre64):

i suggest you tone down your language. and, we are not providing a free answering service, we are offering a way to learn the material.

OpenStudy (amistre64):

my first attempt was to simplify it .... since that did not work i resorted to the longer, more detailed version of it all.

OpenStudy (anonymous):

So the answer is .6875?

OpenStudy (anonymous):

Or 11/16

OpenStudy (amistre64):

11/16 is correct either way is fine, but the format is up to whoever is doing the grading.

OpenStudy (anonymous):

Thanks I just forgot how to do this from the beginning of the year

OpenStudy (amistre64):

you did fine :) good luck